OpenStudy (anonymous):

If the force acting on an object stays the same, then the acceleration of the object is inversely proportional to its mass. If an object with a mass of 9 kilograms accelerates at a rate of 6 meters per second per second by a force, find the rate of acceleration of an object with a mass of 2 kilograms that is pulled by the same force. These are my choices: A. 3 meters per second per second B. 27 meters per second per second C. 18 meters per second per second D. 24 meters per second per second

OpenStudy (welshfella):

inverse proportion A = k/m where k is a constant A = 6 and m = 9: 6 = k / 9 solve for k

OpenStudy (welshfella):

having found the value of k plug it into the formula then plug in m = 2 to work out the value of A

OpenStudy (anonymous):

Thank you so much!